OpenStudy (anonymous):

A quadratic function is a function of the form y=ax2+bx+c where a, b, and c are constants. Given any 3 points in the plane, there is exactly one quadratic function whose graph contains these points. Find the quadratic function whose graph contains the points (0,3), (2,15), and (−1,0).

OpenStudy (raden):

distribute all point into y=ax2+bx+c, we have 3 equations : ((),3) --> 3 = a(0)^2 + b(0) + c, or c = 3 ... (1) (2,15) --> 15 = a(2)^2 + b(2) + c = 15 or 15 = 4a + 2b + c beause c = 3, therefore the equa above can be 15 = 4a + 2b + 3 or 4a + 2b = 12 (divide by 2) 2a + b = 6 ... (2) (-1,0) --> 0 = a(-1)^2 + b(-1) + c 0 = a - b + c again, subtitute c = 3 we get a - b + 3 = 0 a - b = -3 ... (3) by using elimination and subtitution method, solve for a and b for both equations (2 and 3) 2a + b = 6 a - b = -3 (solve for a and b) continue it ...

OpenStudy (anonymous):

ohhh ok. thank you!